OpenStudy (anonymous):

Write the expression in radical form and simplify (a^7b^3)^2/3

OpenStudy (anonymous):

\[\left( a^7b^3 \right)^\frac{2}{3}\]

OpenStudy (anonymous):

distribute the exponents first

OpenStudy (anonymous):

shoot 2/5

OpenStudy (anonymous):

a^14/5b6/5

OpenStudy (anonymous):

\[a^{14/5}b^{6/5}\]

OpenStudy (anonymous):

\[\left( a^7b^3 \right)^\frac{2}{5}\]

OpenStudy (anonymous):

that is what it is supposed to look like

OpenStudy (anonymous):

\[\sqrt[5]{a^{14}b^6}\]

OpenStudy (anonymous):

the 5 outside and the multipleshe inside correct